OpenStudy (anonymous):

hey the width of a rectangle is seven less than half its length the perimeter is 50

OpenStudy (anonymous):

The trick to these kinds of problems is to think in terms of variables instead of words. When you see "the width of a rectangle is..." consider the information that statement gives up: we're dealing with a rectangle, and its width can be expressed in terms of some other variable. We know that the perimeter of a rectangle is given this way: 2L + 2W = P. We're given a lovely expression of W in terms of L: " width is seven less than half the length" which we can relate into an equation\[W=(\frac{ 1 }{ 2 } L) -7\] Now that we only have one variable, it's just a few steps from here to determine L. Let's plug that expression we got for W into the original equation... it should looks like this: \[2L + 2(\frac{1}{2}L -7) = 50\] With me so far? :)

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Sorry, lost my connection for a minute there :\ are you looking for L or W?